Use the Chain Rule to find dz/dt. (Enter your answer only in terms of t.)

z = sin(x + 9y), x = 5t6, y = 3/t

dz/dt = ?

Use the Chain Rule to find dz/dt. (Enter your answer only in terms of t.) z = sin(x + 9), x = 5t6, y = 3/t dz/dt =
Answer #1

Z= sin(x + gy) diff wort t dz - cos(x+) [ * + Poht] © also x=5 to 4 y =3/+ dx = 5x6 + 12 = -3/22 N da = 3045 dt put the val